Use slope intercept form to write the equation of the line: perpendicular to y=(2/5)x+4 and passing through (6,-2)
Answer by mananth(16949)

the line is y= (2/5)x +4 . compare it with y = mx +b where m is the slope.
slope = 2/5
..
The line is perpendicular to this line and passing through ( 6,-2)
the slope of this line will be (-5/2)
Find b in the equation y= mx +b. plug the value (6,-2) in the equation y = mx +b to find b
-2=(-5/2)*6+b
-2=-15+b
b=13
..
The equation of the line will be y=(-5/2)x +13